OpenStudy (anonymous):

Hooke's law. The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 43 cm when the weight is 4 kg, what is the distance when the weight is 5 kg?

OpenStudy (stormfire1):

F= -kx (4kg)(-9.8 m/s^2) = -k(.43m) k = ~91 (5kg)(-9.8 m/s^2) = -91(x) x = .538m = 53.8cm

OpenStudy (anonymous):

so the answer is...?

OpenStudy (stormfire1):

53.8cm

OpenStudy (anonymous):

it says its wrong. It says Type an integer or a simplified fraction.

OpenStudy (stormfire1):

did you try rounding it to 54 cm?

OpenStudy (anonymous):

....no lol

OpenStudy (anonymous):

it said thats wrong to

OpenStudy (stormfire1):

then I have no idea :) I'm pretty sure it's 53.7cm. Maybe someone else will give it a shot and see if they come up with the same thing

OpenStudy (stormfire1):

err... 53.8

OpenStudy (anonymous):

k Thx

OpenStudy (stormfire1):

I see that it accepted 214/4...which is 53.75. That's why I hate doing math online...the systems are too strict with the form of the answer.